# Resisting Quantum Key Distribution Attacks Using Quantum Machine Learning

### v1

@@ III-B Generation
Key_Length QBER Measurement_entropy Signal_detection_rate Decoy_detection_rate Signal_loss_rate Decoy_loss_rate Avg_Photon_time Whole_key_time Arrival_var Arrival_dev Label
253 0.0553 0.2981 0.7311 0.7545 0.2691 0.2455 0.0684 35.34 0.0019 0.0359 normal
247 0.3927 0.9661 0.7510 0.7110 0.2489 0.2889 0.1879 97.13 0.0213 0.1059 mitm_attack
284 0.0739 0.3798 0.7112 0.6446 0.2442 0.3013 0.0923 47.62 0.0059 0.0653 pns_attack
240 0.2542 0.8169 0.7437 0.7679 0.2563 0.2321 0.0932 49.02 0.0071 0.0672 trojan_horse_attack
311 0.0129 0.0990 0.8663 0.8738 0.1337 0.1262 0.0630 44.12 0.0024 0.0380 wavelength_dependent_trojan_attack
454 0.1850 0.6882 0.9090 0.8943 0.0901 0.1057 0.0697 48.79 0.0040 0.0479 rng_attack
142 0.0845 0.4178 0.4340 0.4265 0.2740 2.6567 0.0581 40.69 0.0096 0.0734 detector_blinding_attack
344 0.1744 0.6585 0.8198 0.8815 0.1820 0.1185 0.0653 45.69 0.0025 0.0374 combined_attack
@@
The dataset contains 11 features, but only 9 features have been selected for training the proposed model, excluding the Signal Detection Rate and Decoy Detection Rate. This decision is based on the fact that the Signal Loss Rate and Decoy Loss Rate are already included, and they imply the same information. Removing these features helps prevent overfitting and ensures that the model generalizes well.
@@
where $\mu$ represents the mean of the feature values, and $\sigma$ denotes the standard deviation. Additionally, the categorical labels are encoded using the one-hot encoding technique to ensure that the model correctly interprets the categorical class information.
@@
The dataset is split into training and testing sets with an 80%-20% ratio to ensure sufficient training data while maintaining a reliable evaluation set.
@@
Gaussian noise is added to the training set to improve model robustness and enhance the generalization capability. This noise follows a normal distribution:
@@
where $\mathcal{N}(\mu,\sigma^{2})$ represents Gaussian noise with mean $\mu=0$ and standard deviation $\sigma=0.05$. This ensures that the model is exposed to slightly varied training inputs, making it more resilient to real-world variations.
III-B1 Input Representation
The input batch size is set to 64 samples per batch, ensuring a balance between training efficiency and model stability. The sequence length is fixed at 9 timesteps, corresponding to the 9 selected features extracted from the QKD system.
@@ IV Evaluation
This section discusses the evaluation criteria and presents the performance results of the proposed hybrid QLSTM model compared to classical CNN and LSTM models, followed by a comparative analysis with existing work in the literature. All models were evaluated using the dataset discussed in Section III-B, after applying the same preprocessing steps detailed in the same section, and trained using the same setup described in Section IV-A to ensure fair evaluation.
@@ IV-A Setup
The three architectures QLSTM, LSTM, and CNN were trained under identical conditions. Training was conducted on a GPU (Google Colab); both model parameters and data tensors were moved to the GPU device to leverage accelerated matrix operations and, in the case of the QLSTM, faster quantum‐circuit simulations. We used PyTorch’s CrossEntropyLoss, which applies a softmax over the logits followed by the multi‐class log‐loss. Optimization was performed using the AdamW optimizer with initial learning rate $5\times 10^{-4}$ and weight decay $1\times 10^{-4}$. A Cosine Annealing Warm Restarts scheduler with restart period $T_{0}=50$ epochs was applied to cyclically modulate the learning rate. Mini‐batches of 64 samples were used throughout the training. Early stopping with patience of 5 epochs monitored the validation loss, halting training if no improvement was observed for five consecutive epochs and thus avoiding overfitting. Each model was trained for 10, 20, and 50 epochs under this same protocol. These parameters are summarized in Table VII.
@@
Scheduler
Cosine Annealing
Warm Restarts ($T_{0}=50$)
@@ IV-B Metrics
Represents the overall correctness of the model by measuring the proportion of all predictions (both attack and normal) that are correctly classified.
@@
Measures how many predicted attacks are actual attacks, reducing false positives.
@@
Evaluates how well the model detects actual attacks, minimizing false negatives.
@@
Balances both precision and recall, useful when there is an imbalance between normal and attack labels.
@@
Furthermore, the model’s performance is compared against classical CNN and LSTM models, all trained under the same conditions, to assess the proposed model’s effectiveness on multi-class classification. Although CNN model is traditionally designed for spatial data, in this study, we adapted it using 1D convolutional layers by reshaping the input such that the 9 features form a one-dimensional structure.
@@ V Results
The evaluation results for all models demonstrate that the proposed QLSTM model outperforms both LSTM and CNN in detecting quantum attacks on the QKD system. Table VIII shows that even in the early stages of training (10 epochs), QLSTM achieves an accuracy of 88.3%, already higher than LSTM (85.8%) and CNN (86.8%). This early lead reflects QLSTM’s ability to capture temporal and statistical patterns present in quantum attack traffic, which are more challenging for classical models to extract. At 20 epochs, QLSTM continues to improve, reaching 93.8% accuracy, and by 50 epochs it achieves 94.7% accuracy, 95.1% precision, and 94.7% F1-score—surpassing both classical models across all metrics. While LSTM shows a steady but limited improvement, it peaks at 88.2% accuracy before dropping at 50 epochs, suggesting overfitting. CNN shows a similar trend, slightly improving until 20 epochs but then decreasing in performance, indicating limited ability to generalize across different attack scenarios. Figure 3 (a) illustrates how the accuracy evolves as the number of epochs increases for QLSTM, LSTM, and CNN. Unlike the classical models, which tend to plateau or degrade with more training, QLSTM continues to improve, reflecting its deeper learning capacity. Figures 3 (b, c, and d) further confirm QLSTM’s superior performance, showing that it consistently achieves the highest precision, recall, and F1-score. While CNN initially maintains balanced metrics, it begins to overfit, particularly at higher epochs. LSTM retains moderate balance but lacks the expressiveness to fully distinguish between quantum attack types and normal traffic in QKD processes.
@@
Figure 3: Performance comparison of different models: a) Accuracy vs. Epochs for Hybrid QLSTM, LSTM, and CNN models, b) Precision vs. Epochs for Hybrid QLSTM and Classical LSTM, c) Recall vs. Epochs for Hybrid QLSTM and Classical LSTM, d) F1-score vs. Epochs for Hybrid QLSTM and Classical LSTM, and e) Loss vs. Epochs for Hybrid QLSTM, LSTM, CNN models.
@@
Figure 4: Confusion matrices for the evaluated models: (a) CNN, (b) LSTM, and (c) Hybrid QLSTM.
@@
To further analyze model behavior, Figure 4 presents the confusion matrices for CNN, LSTM, and QLSTM at 50 epochs. QLSTM demonstrates the most consistent and accurate classification across all eight classes. It successfully detects Combined Attack (class 0), Detector Blinding Attack (class 1), Intercept and Resend Attack (class 2), Normal (class 3), PNS Attack (class 4), RNG Attack (class 5), Trojan Horse Attack (class 6), and Wavelength Dependent Trojan Attack (class 7). Most misclassifications in QLSTM are limited to overlaps between Normal and PNS Attack, which share similar photon-level patterns. In comparison, LSTM and CNN exhibit broader confusion, especially with Detector Blinding and Trojan Horse attacks. CNN also struggles more with differentiating Normal and PNS Attack, showing higher false positives and negatives. While LSTM handles Detector Blinding Attack slightly better, it still misclassifies Intercept and Resend and Normal traffic due to their overlapping behaviors. These results confirm that QLSTM is more capable of learning the complex quantum patterns needed for accurate classification in QKD environments.
@@
Table VIII presents the final evaluation metrics after 50 epochs of training. The QLSTM achieves better performance across all considered metrics, confirming its enhanced ability to detect and classify attacks in the QKD system.
@@ VI Conclusion
In this work, we addressed security vulnerabilities in QKD by developing a hybrid intrusion detection system based on QLSTM. While QKD provides theoretically unbreakable encryption, its real-world implementations remain vulnerable to various attacks, such as Intercept-Resend, PNS, and Trojan Horse attacks. To enhance the security of QKD networks, this work integrates quantum-enhanced learning into a deep learning-based intrusion detection framework.
A key contribution of this work is the development of a realistic QKD dataset due to the lack of publicly available datasets for QKD attack detection. The dataset includes essential quantum security metrics, such as QBER, measurement entropy, signal and decoy loss rates, and time-based metrics, ensuring an accurate representation of practical QKD conditions. The dataset incorporates three attack scenarios: intercept-and-resend, PNS, Trojan Horse, wavelength-dependent Trojan Horse, RNG, Detector Blinding, and combined attacks.
This work introduces a QLSTM model for detecting attacks on QKD systems by capturing the temporal patterns in quantum metrics. The model’s performance was evaluated against classical deep learning models, including LSTM, and CNN. The results showed that QLSTM not only achieved high accuracy of 93.7% after 50 training epochs on the proposed dataset, but also demonstrated stronger generalization capabilities. Overall, the results demonstrate the effectiveness of the proposed QLSTM model for intrusion detection in QKD systems. This promising performance suggests that hybrid quantum-classical approaches could play an important role in enhancing the security of future QKD networks.
Despite the promising results, the proposed model has certain limitations, which provide several areas for future investigations. First, the current training process utilized fixed hyperparameters, leaving room for further exploration through fine-tuning. Adjusting the learning rate, batch size, and optimizer settings could enhance the model’s convergence and generalization. Second, the current evaluation focused on normal QKD and three attack scenarios. Considering more attack scenarios within the dataset would enable a more comprehensive evaluation of the model’s robustness. Third, the current dataset simulates a 20-bit key, so future work should aim to generate datasets for 256-bit keys to better reflect real-world cryptographic requirements. Fourth, adopting alternative or more advanced models, such as Gated Recurrent Units (GRUs), or quantum-enhanced GRUs, could offer improvements in both performance and computational efficiency. Lastly, benchmarking against proposed models from the literature would provide further insights into the strengths and weaknesses of the QLSTM model in real-world applications.

### v2

@@ III-D Generation
Sifted_Key_Length QBER Measurement_entropy Signal_detection_rate Decoy_detection_rate Avg_Photon_time Whole_key_time Arrival_var Arrival_dev Label
253 0.0553 0.2981 0.7311 0.7545 0.0684 35.34 0.0019 0.0359 normal
247 0.3927 0.9661 0.7510 0.7110 0.1879 97.13 0.0213 0.1059 mitm_attack
284 0.0739 0.3798 0.7112 0.6446 0.0923 47.62 0.0059 0.0653 pns_attack
240 0.2542 0.8169 0.7437 0.7679 0.0932 49.02 0.0071 0.0672 trojan_horse_attack
311 0.0129 0.0990 0.8663 0.8738 0.0630 44.12 0.0024 0.0380 wavelength_dependent_trojan_attack
454 0.1850 0.6882 0.9090 0.8943 0.0697 48.79 0.0040 0.0479 rng_attack
142 0.0845 0.4178 0.4340 0.4265 0.0581 40.69 0.0096 0.0734 detector_blinding_attack
344 0.1744 0.6585 0.8198 0.8815 0.0653 45.69 0.0025 0.0374 combined_attack
@@
The dataset contains 11 features, but only 9 are used to train the proposed model, excluding the Signal Detection Rate and Decoy Detection Rate. This decision is based on the fact that the Signal Loss Rate and the Decoy Loss Rate are already included and convey the same information. Removing these features helps prevent overfitting and ensures the model generalizes well.
@@
where $\mu$ represents the mean of the feature values, and $\sigma$ denotes the standard deviation. In addition, categorical labels are encoded as one-hot vectors to ensure that the model correctly interprets class information.
@@
The dataset is divided into training and test sets at 80%-20% to ensure sufficient training data while maintaining a reliable evaluation set.
@@
Gaussian noise is added to the training set to improve model robustness and enhance generalization. This noise follows a normal distribution:
@@
where $\mathcal{N}(\mu,\sigma^{2})$ represents Gaussian noise with mean $\mu=0$ and standard deviation $\sigma=0.05$. This ensures that the model is exposed to slightly varied training input, making it more resilient to real-world variations.
III-D1 Input Representation
The input batch size is set to 64 samples per batch, balancing training efficiency and model stability. The sequence length is fixed at 9 time steps, corresponding to the 9 selected features extracted from the QKD system.
III-D2 Physical Models and Dataset Rationality.
The dataset was constructed according to the physical parameters specified in the IEEE 2021[25] standards for BB84 systems with decoy-state. Specifically, signal and decoy pulse intensities ($\mu_{s}$, $\mu_{d}$) and detection probabilities were embedded within the dataset generation process, eliminating the need to list them explicitly as independent variables. Each data record implicitly reflects the photon statistics, pulse repetition rate, and total number of transmitted pulses $N_{total}$ used to calculate the Quantum Bit Error Rate (QBER) through
$\text{QBER}=\frac{N_{error}}{N_{signal}+N_{decoy}}.$ (11)
By encoding the physical quantities into the signal and decoy detection rates, we ensure that the dataset preserves the essential statistical behavior of photon-level measurements. This formulation provides a physically meaningful representation of the quantum communication features, enabling a faithful yet simulation-based reproduction of the QKD channel statistics without direct access to optical hardware.
@@ IV Evaluation
This section discusses the evaluation criteria and presents the performance results of the proposed hybrid QLSTM model compared to classical CNN and LSTM models, as well as additional baseline architectures including an Artificial Neural Network (ANN) and a Recurrent Neural Network (RNN), followed by a comparative analysis with existing work in the literature. All models were evaluated using the dataset discussed in Section III-D, after applying the same preprocessing steps detailed in the same section, and trained using the same setup described in Section IV-A to ensure fair evaluation.
@@ IV-A Setup
The three architectures QLSTM, LSTM, CNN, ANN, and RNN were trained under identical conditions. Training was conducted on a GPU (Google Colab); both the model parameters and the data tensors were moved to the GPU device to use accelerated matrix operations and, in the case of QLSTM, faster quantum‐circuit simulations. We used PyTorch’s CrossEntropyLoss, which applies a softmax on the logits followed by the multi‐class log‐loss. Optimization was performed using the AdamW optimizer with initial learning rate $5\times 10^{-4}$ and weight decay $1\times 10^{-4}$. A Cosine Annealing Warm restart scheduler with period $T_{0}=50$ epochs was applied to cyclically modulate the learning rate. Mini‐batches of 64 samples were used throughout the training. Early stopping with a patience of 5 epochs monitored the validation loss, halting training if no improvement was observed for 5 consecutive epochs, thereby avoiding overfitting. Each model was trained for 10, 20, and 50 epochs under this same protocol. These parameters are summarized in Table VII.
@@
Scheduler Cosine Annealing Warm Restarts ($T_{0}=50$)
@@ IV-B Metrics
It represents the overall correctness of the model by measuring the proportion of all predictions (both attack and normal) that are correctly classified.
@@
It measures how many predicted attacks are actually executed, reducing false positives.
@@
It evaluates how well the model detects actual attacks, minimizing false negatives.
@@
It balances both precision and recall, which is useful when there is an imbalance between normal and attack labels.
@@
Furthermore, the model’s performance is compared with those of classical CNN, LSTM, ANN, and RNN models, all trained under the same conditions, to assess the proposed model’s effectiveness for multi-class classification. Although the CNN model is traditionally designed for spatial data, we adapted it for 1D convolutional layers by reshaping the input so that the 9 features form a 1D structure.
@@ V Results
The evaluation results for all models demonstrate that the proposed QLSTM model outperforms both LSTM and CNN in detecting quantum attacks on the QKD system, along with the additional ANN, Random Forest and RNN models introduced for a broader comparison. Table VIII shows that even in the early stages of training (10 epochs), QLSTM achieves an accuracy of 88.3%, already higher than LSTM (85.8%), Random Forest (82.9%), CNN (86.8%), and ANN (89.2%), while significantly surpassing RNN (67.9%). This early lead reflects QLSTM’s ability to capture temporal and statistical patterns present in quantum attack traffic, which are more challenging for classical models to extract. At 20 epochs, QLSTM continues to improve, reaching 93.8% accuracy, and by 50 epochs it achieves 94.7% accuracy, 95.1% precision and 94.7% F1-score—surpassing both classical models across all metrics. Although LSTM shows a steady but limited improvement, it peaks at 88.2% accuracy before dropping at 50 epochs, suggesting overfitting. CNN shows a similar trend: it improves slightly until 20 epochs, then performance decreases, indicating limited ability to generalize across different attack scenarios. ANN follows a similar trend, achieving consistent but lower improvements compared to QLSTM, while RNN shows the weakest overall performance, struggling to generalize across quantum attack classes due to its limited temporal representation capacity. Figure 3 (a) illustrates how the accuracy evolves as the number of epochs increases for QLSTM, LSTM ,Random Forest,and CNN. Unlike classical models, which tend to plateau or degrade with more training, QLSTM continues to improve, reflecting its deeper learning capacity. Figures 3 (b, c, d and f) further confirm the superior performance of QLSTM, showing that it consistently achieves the highest precision, recall, and F1-score. While CNN initially maintains balanced metrics, it begins to overfit, particularly at higher epochs. Random Forest establishes a reliable baseline with high stability, yet it hits a distinguishable performance ceiling, failing to capture the complex sequential dependencies of quantum attacks as effectively as the advanced architectures. LSTM maintains a moderate balance but lacks the expressiveness to fully distinguish between quantum attack types and normal traffic in QKD processes, whereas ANN remains more robust but still underperforms QLSTM, and RNN exhibits noticeable instability across metrics.
@@
Figure 3: Performance comparison of different models: Hybrid QLSTM, LSTM, CNN, RNN, ANN, and Random Forest: a) Accuracy vs. Epochs, b) Precision vs. Epochs, c) Recall vs. Epochs, d) F1-score vs. Epochs, e) Loss vs. Epochs, and f) Final Accuracy Comparison.
@@
Figure 4: Confusion matrices for the evaluated models: (a) CNN, (b) LSTM, (c) Hybrid QLSTM, (d) ANN, (e) RNN, and (f) Random Forest.
@@
To further analyze model behavior, Figure 4 presents the confusion matrices for ANN, RNN, CNN, LSTM, and QLSTM at 50 epochs. QLSTM demonstrates the most consistent and accurate classification in all eight classes. It successfully detects Combined Attack (class 0), Detector Blinding Attack (class 1), Intercept and Resend Attack (class 2), Normal (class 3), PNS Attack (class 4), RNG Attack (class 5), Trojan Horse Attack (class 6), and Wavelength Dependent Trojan Attack (class 7). Most misclassifications in QLSTM are limited to overlaps between Normal and PNS Attack, which share similar photon-level patterns. This overlap arises from their nearly identical photon-level statistics—both classes exhibit similar signal and decoy photon distributions, making the distinction subtler for any model. Adjusting the attack probabilities or photon intensity parameters during dataset generation could help reshape these class boundaries, leading to clearer separability in future experiments. In comparison, LSTM and CNN exhibit broader confusion, especially with Detector Blinding and Trojan Horse attacks. CNN also struggles more to differentiate Normal and PNS Attack, resulting in higher false positives and negatives. Although LSTM handles the Detector Blinding Attack slightly better, it still misclassifies Intercept, Resend, and Normal traffic due to their overlapping behaviors. The ANN model shows relatively stable results, performing comparably to CNN but with a slightly improved distinction between Normal and PNS Attack classes. However, it still exhibits limited confusion across high-similarity classes, reflecting its inability to fully capture quantum-feature interactions. Conversely, RNN exhibits the weakest performance among all models, with pronounced confusion across most attack classes—particularly between Combined, Intercept, and Resend, and RNG Attacks—indicating that its sequential dependencies alone are insufficient to model the quantum correlations effectively.
@@
These results confirm that QLSTM outperforms other methods in learning complex quantum patterns required for accurate classification in QKD environments. Table VIII presents the final evaluation metrics after 50 epochs of training. The QLSTM achieves better performance across all metrics considered, confirming its enhanced ability to detect and classify attacks in the QKD system.
The observed performance aligns with the theoretical expectations discussed in Section II. Hybrid QLSTM demonstrated faster convergence, higher accuracy, and stronger generalization compared to classical models, confirming the advantage of quantum-enhanced feature representation in sequential learning. These results suggest that the variational quantum circuits within the QLSTM effectively captured temporal dependencies and nonlinear correlations in the QKD parameters, leading to more stable optimization and improved discrimination between attack classes. In addition, the QLSTM’s resistance to overfitting at later epochs indicates that quantum properties such as superposition and entanglement introduce an implicit regularization effect, thereby enhancing learning efficiency. Overall, these findings validate that integrating quantum computation into recurrent architectures provides measurable benefits for complex temporal modeling in quantum key distribution security.
@@
To complement the performance comparison across ANN, CNN, RNN, LSTM, and the proposed Hybrid QLSTM, we also benchmarked the computational requirements of each model under identical hardware and training settings. All experiments were executed using the same batch size, optimizer, learning rate, early-stopping configuration, and GPU (NVIDIA A100). Table IX reports the number of trainable parameters, average wall-clock time per epoch, total training time for 50 epochs, inference latency per sample, and maximum GPU memory usage for each architecture.
TABLE IX: Computational Resource and Training-Time Benchmarking Across Models
Model Params Time/Epoch (s) 50-Epoch Time (s) Infer/Sample (ms) GPU Mem (MB)
ANN $1.4\times 10^{4}$ 0.10 5.0 0.02 120
RNN $1.1\times 10^{4}$ 0.09 4.5 0.02 100
CNN $2.1\times 10^{4}$ 0.18 9.0 0.03 150
LSTM $4.3\times 10^{4}$ 0.30 15.0 0.05 190
Random Forest $5.1\times 10^{5}$ - 8.9 0.07 0
QLSTM $3.9\times 10^{3}+72$ 0.41 20.5 0.06 225
As shown in Table IX, the Hybrid QLSTM exhibits the highest time per epoch (0.41 s) and the largest overall training time (20.5 s for 50 epochs). This overhead is expected, as each forward and backward pass requires evaluating variational quantum circuits, which involve parameterized unitary transformations and repeated sampling. Despite this added cost, the Hybrid QLSTM maintains a notably lower number of classical parameters than LSTM and CNN and achieves the best accuracy, precision, recall, and F1-score among all models (Table VIII). These results indicate that QLSTM offers a favorable trade-off: a modest increase in computational cost in exchange for significantly improved detection performance in QKD intrusion-detection scenarios.
@@ VI Conclusion
This work addressed the security vulnerabilities that persist in practical QKD implementations by introducing a hybrid quantum–classical intrusion detection system based on a QLSTM architecture. Although QKD offers theoretically unbreakable security at the protocol level, its physical implementations remain vulnerable to a variety of sophisticated attacks—including Intercept–Resend, PNS, Trojan-Horse, and wavelength-dependent Trojan-Horse attacks, RNG tampering, detector blinding, and composite strategies. To mitigate these threats, our approach integrates quantum-enhanced temporal learning within a deep learning framework to strengthen the resilience of QKD networks. A key contribution of this work is the construction of a realistic and comprehensive QKD security dataset, which addresses the lack of publicly available resources for QKD intrusion detection. The dataset incorporates essential quantum communication and security parameters—QBER, entropy metrics, decoy-state statistics, photon count distributions, and timing-based indicators—to emulate real-world QKD operational behavior. Multiple attack scenarios were included to ensure rigorous and diverse model evaluation. Using this dataset, we developed a QLSTM-based IDS capable of capturing temporal, statistical, and quantum-derived signatures for accurate detection of abnormal behavior in QKD traffic. Benchmarking against strong classical baselines—including LSTM and CNN models—demonstrated that the proposed QLSTM achieves a detection accuracy of 94.7% after 50 epochs and exhibits superior generalization capabilities. Although the Random Forest model established a robust baseline with high stability and computational efficiency (achieving $\sim$82.6% accuracy), it reached a distinguishable performance ceiling, confirming that capturing the complex sequential dependencies of sophisticated quantum attacks requires the advanced temporal expressiveness found in quantum-enhanced architectures. These findings highlight the growing potential of hybrid quantum–classical models to improve security monitoring in emerging quantum communication infrastructures.
Despite these promising results, several limitations remain. First, the current training pipeline uses fixed hyperparameters; more systematic hyperparameter optimization could further enhance model stability and generalization. Second, although the dataset includes eight representative attack types, the attack space in practical QKD networks is broader. Extending the dataset with additional attack models would support a more comprehensive robustness analysis. Third, the dataset used in this study simulates semi-realistic QKD behavior via simplified, simulation-friendly variants of known attacks. For example, the implemented PNS attack omits optical-loss compensation steps, and the Trojan-Horse variants employ non-calibrated power settings and wavelength models rather than full hardware-level emulation. Fourth, although algorithmic simulations of the QLSTM framework demonstrate strong performance, current Noisy Intermediate-Scale Quantum (NISQ) hardware lacks the fidelity, qubit connectivity, and coherence times required for accurate runtime benchmarking. As quantum processors mature, evaluating QLSTM on real devices will help quantify the true computational overhead. Fifth, exploring alternative architectures—such as GRUs, quantum-enhanced GRUs, hybrid encoder–decoder models, or attention-based QML variants—may yield improvements in accuracy and efficiency. Sixth, comparing QLSTM performance with recent state-of-the-art QKD intrusion detection methods will provide deeper insights into its relative strengths and limitations. Lastly, future work will incorporate quantitative evaluation of Trojan-Horse effects by computing parameters such as $(p_{\text{inj}},\rho_{\text{ret}},F_{\text{Eve}})$ through optical simulations or controlled laboratory experiments. This will enable direct validation of the proposed min-entropy and leakage bounds and further enhance the QLSTM’s sensitivity to side-channel deviations in the physical layer. Furthermore, integrating calibrated flux control, adaptive loss modeling, and hardware-level optical parameters to narrow the gap between simulated and experimental conditions.